Let the angle be the angle that the vector makes with

the measured counterclockwise from that axis. Find

the angle for a vector that has the following components: 

(a)Ax=2.50m, Ay=-0.500m


1
Expert's answer
2022-02-01T15:34:03-0500

Given:

"A_x=2.50\:m, \;A_y=-0.500\:m"


The angle that the vector makes with "x"-axis

"\theta=\tan^{-1}\frac{A_y}{A_x}=\tan^{-1}\frac{-0.500}{2.50}=349^\circ"